how do i figure out -7x-3x+2=8x-8?
Answer by checkley79(3341)

-7x-3x+2=8x-8
-7X-3X-8X=-8-2
-18X=-10
X=-10/-18
X=5/9 ANS.
PROOF:
-7*(5/9)-3*(5/9)+2=8*(5/9)-8
-35/9-15/9+2=40/9-8
-50/9-40/9=-8-2
-90/9=-10
-10=-10